OpenStudy (anonymous):

evaluate 10 plus(-3)-(-4)-(positive6)

OpenStudy (anonymous):

10 - 3 + 4 - 6 = 14 - 9 = 5

OpenStudy (anonymous):

?=10+(-3)-(-4)-6 ?=7-(-4)-6 ?=11-6 ?=5

OpenStudy (anonymous):

double negative = positive so instead of doing -3-(-4) you can do -3+4 I find it easier :)

OpenStudy (anonymous):

ty